# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ A.1 Coherent displacement of driven cavity mode
$\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}$
$-2i\Omega(t)\cos(\omega_{d}t),$ (37)
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(a+a)^{{\dagger}}\sigma_{j}^{z}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (38)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})-2ig_{j}(a^{{\dagger}}+a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}(t)$ corresponds to the input noise field of the cavity mode, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$. A coherent driving on the cavity induces a coherent state component on this mode, and thus it is convenient to displace it using the transformation $a\rightarrow\alpha+\delta a$ as in Eq. (7). Here, $\delta a$ corresponds to the quantum fluctuation of the cavity mode around its classical value $\alpha$. By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation reads
$\delta\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]\delta a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}(t),$ (40)
@@
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(\delta a+\delta a^{{\dagger}})\sigma_{j}^{z}$
$+2ig_{j}{\rm Re}\{\alpha\}\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}$
$-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (41)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-2ig_{j}(\delta a^{{\dagger}}+\delta a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-4ig_{j}{\rm Re}\{\alpha\}(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
@@
Here, $\alpha(t)$ is determined via the classical differential equation (8).
The master equation associated to the above Langevin equations (40)-(42) with the displaced cavity reads
$\dot{\rho}^{\prime}(t)={}$ $-i[H^{\prime}(t),\rho^{\prime}]+\kappa{\cal{D}}[\delta a]\rho^{\prime}+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}$
@@
$H^{\prime}(t)={}$ $\omega_{c}\delta a^{{\dagger}}\delta a+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
$+\sum_{j=1}^{N}2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
Notice that the master equation (43) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (44), which acts on the qubits rather than on the cavity as indicated in Eq. (10). This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
The displaced master equation (43) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$ because most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is only weakly populated. This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits. In addition, if we consider the cavity in the weak coupling and bad cavity limit given by Eq. (11), the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq. (44). On the one hand, since $g_{j}\ll\omega_{c}+\omega_{q}^{j}$ and the fluctuation is weakly populated, we can approximate the qubit-fluctuation coupling as
$g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})\approx g_{j}(\delta a^{\dagger}\sigma_{-}+{\rm h.c.}).$ (45)
@@
$2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})\approx g_{j}|\alpha_{\rm ss}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi)}+{\rm h.c.}).$ (46)
Here, $|\alpha_{\rm ss}|=\Omega_{0}/\sqrt{(\kappa/2)^{2}+\Delta^{2}}$ and $\phi={\rm Arctan}(2\Delta/\kappa)-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile. To demonstrate this last approximation (46), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha(0)=0$. Then, the solution of the classical equation (8) for $\alpha(t)$ reads
$\alpha(t)={}$ $|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}(1-e^{-(\kappa/2-i\Delta)t})$
$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa/2+2i\omega_{d})t}),$ (47)
where we have also used the inequalities $|\Delta|,\kappa\ll\omega_{d}$ valid for our parameter conditions. The first term in (47) reaches the steady state value $\sim|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ in a timescale $\sim 1/\kappa$ and the second term causes the fast but low amplitude oscillation observed in Fig. 2(a). Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss}|\gg 1$ as it is the case for our parameters. Since in bad cavity limit (11) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa$, we can use the steady state value $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits. Similarly, after the cavity reaches the steady state and it is switched off ($f(t)=0$), the dynamics in Eq. (8) predicts an exponential decay of $\alpha(t)$ with rate $\sim\kappa$. In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha(t)\approx 0$ for $f(t)=0$.
In summary, under the above approximations we have $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}f(t)$ so that when replacing it in the last term of Eq. (44), and applying RWA provided $g_{j}|\alpha_{\rm ss}|\ll\omega_{d}$, we obtain Eq. (46). Replacing Eqs. (45)-(46) in (44), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (43) with the RWA Hamiltonian:
$H^{\prime}(t)\approx{}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
@@
Here, we have defined $\Omega_{\rm cm}^{j}=g_{j}|\alpha_{\rm ss}|$ as in Eq. (14) and absorbed the constant phase $\phi$ in the definition of the qubit and fluctuation operators. The qubit detunings read $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ as in the main text.
A.2 Adiabatic elimination of cavity fluctuations
As discussed in the previous subsection, in the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state, with its quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$. In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only. To do so, we formally integrate Eq. (40) and apply Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$. As a result, we get
$\delta a(t)\approx{}$ $-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t)$
${\color[rgb]{0,0,0}-i\sum_{j}\frac{g_{j}}{\kappa/2+i(2\omega_{d}-\Delta)}\sigma_{j}^{+}(t).}$ (49)
Replacing the above expressions in the original quantum Langevin equations (38)-(39), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations
@@ A.2 Adiabatic elimination of cavity fluctuations
$+i\Omega_{\rm cm}^{j}f(t)\sigma_{j}^{z}+\sum_{l\neq j}\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{z}\sigma_{j}^{-}+\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{z}a_{\rm in}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$
@@
$-2\sum_{l\neq j}\left[\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{+}\sigma_{l}^{-}+{\rm h.c.}\right]$
$-2\sqrt{\gamma^{j}_{\rm cm}}\left[\sigma_{j}^{+}a_{\rm in}+{\rm h.c.}\right]-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]$
$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (51)
Here, the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (14)-(17).
Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs. (50)-(51). This effective master equation, which incorporates the adiabatic elimination of the cavity mode, is given in Eqs. (12)-(13).

### v3

@@ A.2 Coherent displacement of driven cavity modes
$\dot{a}_{m}={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]a_{m}-i\sum_{j=1}^{N}g_{j,m}(\sigma_{j}^{-}+\sigma_{j}^{+})$ (44)
$-2i\Omega_{m}(t)\cos(\omega_{d}t)-\sqrt{\kappa_{m}}a_{\rm in}^{m},$
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}$ (45)
$+i\sum_{m}g_{j,m}(a_{m}+a_{m}^{{\dagger}})\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}$
$+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (46)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})$ (47)
$-2i\sum_{m}g_{j,m}(a_{m}^{{\dagger}}+a_{m})(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$
Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}^{m}(t)$ corresponds to the input noise field of the cavity mode $m$, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$. A coherent driving on the cavity induces a coherent state component on this mode, and thus it is convenient to displace it using the transformation $a_{m}\rightarrow\alpha_{m}+\delta a_{m}$ as in Eq. (8). Here, $\delta a_{m}$ corresponds to the quantum fluctuation of the cavity mode around its classical value $\alpha_{m}$. By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation of cavity mode $m$ reads,
$\delta\dot{a}_{m}={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]\delta a_{m}-i\sum_{j=1}^{N}g_{j,m}(\sigma_{j}^{-}+\sigma_{j}^{+})$
$-\sqrt{\kappa_{m}}a_{\rm in}^{m}(t),$ (48)
@@
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+2i\sum_{m}g_{j,m}{\rm Re}\{\alpha_{m}\}\sigma_{j}^{z}$
$+i\sum_{m}g_{j,m}(\delta a_{m}+\delta a_{m}^{{\dagger}})\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}$
$+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (49)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-4i\sum_{m}g_{j,m}{\rm Re}\{\alpha_{m}\}(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-2i\sum_{m}g_{j,m}(\delta a_{m}^{{\dagger}}+\delta a_{m})(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
@@
Here, $\alpha_{m}(t)$ is determined via the classical differential equation,
$\dot{\alpha}_{m}(t)={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]\alpha_{m}(t)-2i\Omega_{0}f(t)\cos(\omega_{d}t),$ (51)
which is a multi-mode generalization of Eq. (10). The master equation associated to the above Langevin equations (48)-(50) with the displaced cavity reads
$\dot{\rho}^{\prime}(t)={}$ $-i[H^{\prime}(t),\rho^{\prime}]+\sum_{m}\kappa_{m}{\cal{D}}[\delta a_{m}]\rho^{\prime}+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}$
@@
$H^{\prime}(t)={}$ $\sum_{m}\omega_{c}^{m}\delta a_{m}^{{\dagger}}\delta a_{m}+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}$
$+\sum_{j=1}^{N}\sum_{m}g_{j,m}(\delta a_{m}+\delta a_{m}^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
$+\sum_{j=1}^{N}\sum_{m}2g_{j,m}{\rm Re}\{\alpha_{m}(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (53)
Notice that, the displaced master equation (52) has the same shape as Eq. (7) except for the inclusion of multi-mode nature and is expressed in terms of the fluctuations $\delta a_{m}$. The driving term in displaced Hamiltonian (53) acts on the qubits rather than on the cavity modes and this is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
Equation (52) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega_{0}|\gg\kappa_{m},|\omega_{c}^{m}-\omega_{d}|$ because most of the cavity photons can be taken into account by the classical field $\alpha_{m}(t)$, and the fluctuations are only weakly populated, $\langle\delta a_{m}^{\dagger}\delta a_{m}\rangle\ll 1$. The steady state solution of Eq. (51) with $f(t)=1$ is given by $\alpha_{\rm ss,m}=-i\Omega_{0}/(\kappa_{m}/2-i\Delta_{m})$, with $\Delta_{m}=\omega_{d}-\omega_{c}^{m}$ the detuning of the $m$-th mode with respect to the cavity drive. This implies that only a few cavity modes with $\omega_{c}^{m}\sim\omega_{d}$ have a significant displacement $\alpha_{m}$ during evolution. In addition, we consider that all cavity modes are weakly coupled to the emitters, which are nearly resonant to the cavity drive, i.e. $g_{j,m}\ll\omega_{q}^{j}\approx\omega_{d}$, so the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq. (53). On the one hand, we can approximate the qubit-fluctuation coupling as
$\sum_{m}g_{j,m}$ $(\delta a_{m}+\delta a_{m}^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
$\approx\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}(\delta a_{m}^{\dagger}\sigma_{j}^{-}+{\rm h.c.}).$ (54)
@@
$\sum_{m}$ $g_{j,m}2{\rm Re}\{\alpha_{m}(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})$ (55)
$\approx\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}|\alpha_{\rm ss,m}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi_{m})}+{\rm h.c.}),$
where $|\alpha_{\rm ss,m}|=\Omega_{0}/\sqrt{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}$ and $\phi_{m}={\rm Arctan}(2\Delta_{m}/\kappa_{m})-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile. To demonstrate this last approximation (55), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha_{m}(0)=0$. Then, the solution of the classical equation (9) for $\alpha_{m}(t)$ reads
$\alpha_{m}(t)={}$ $|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}(1-e^{-(\kappa_{m}/2-i\Delta_{m})t})$
$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa_{m}/2+2i\omega_{d})t}),$ (56)
where we have also used the inequalities $|\Delta_{m}|,\kappa_{m}\ll\omega_{d}$ valid for our parameter conditions. The first term in (56) reaches the steady state value $\sim|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}$ in a timescale $\sim 1/\kappa_{m}$ and the second term causes the fast but low amplitude oscillation observed in Fig. 2(a). Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss,m}|\gg 1$ as it is the case for our parameters. Since in bad cavity limit (12) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa_{m}$, we can use the steady state value $\alpha_{m}(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi_{m})}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits. Similarly, after the cavity reaches the steady state and it is switched off ($f(t)=0$), the dynamics in Eq. (9) predicts an exponential decay of $\alpha_{m}(t)$ with rate $\sim\kappa_{m}$. In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha_{m}(t)\approx 0$ for $f(t)=0$. In summary, under the above approximations we have $\alpha_{m}(t)\approx|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}f(t)$ so that when replacing it in the last term of Eq. (53), and applying RWA provided $g_{j,m}|\alpha_{\rm ss,m}|\ll\omega_{c}^{m},\omega_{q}^{j}\approx\omega_{d}$, we obtain Eq. (55). Replacing Eqs. (54)-(55) in (53), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (52) with the RWA Hamiltonian,
$H^{\prime}(t)\approx{}$ $-\sum_{\omega_{c}^{m}\sim\omega_{d}}\Delta_{m}\delta a_{m}^{{\dagger}}\delta a_{m}-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}$
$+\sum_{j=1}^{N}\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}(\delta a_{m}^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a_{m})$
@@
Here, $\Omega_{\rm cm}^{j}=\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}|\alpha_{\rm ss,m}|$ is the multi-mode generalization of the cavity-mediated driving in Eq. (15) and we have also absorbed the constant phase $\phi_{m}$ in the definition of the qubit and fluctuation operators. The qubit detunings read $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ as in the main text.
A.3 Adiabatic elimination of cavity fluctuations
As discussed in the previous subsection, in the bad cavity limit $g_{j,m},|\delta_{j}|\ll\kappa_{m}$, the cavity modes reach very quickly a coherent steady state, with its quantum fluctuations $\delta a_{m}$ close to the vacuum state $|0\rangle$. In this situation, it is convenient to adiabatically eliminate the cavity fluctuation and obtain effective dynamics for the degrees of freedom of the qubits only. To do so, we formally integrate Eq. (48) and apply Markov approximation provided the qubits’ dynamics evolve slowly on the time-scale $\sim 1/\kappa_{m}$. As a result, we get
$\delta a_{m}(t)\approx{}$ $-i\sum_{j}\frac{g_{j,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa_{m}}a_{\rm in}^{m}(t)}{(\kappa_{m}/2-i\Delta_{m})}$
$-i\sum_{j}\frac{g_{j,m}}{\kappa_{m}/2+i(2\omega_{d}-\Delta_{m})}\sigma_{j}^{+}(t).$ (58)
Replacing the above expressions in the original quantum Langevin equations (45)-(47), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations,
@@ A.3 Adiabatic elimination of cavity fluctuations
$+i\Omega_{\rm cm}^{j}f(t)\sigma_{j}^{z}+\sum_{l\neq j}\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{z}\sigma_{l}^{-}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-}$
$+\sum_{\omega_{c}^{m}\sim\omega_{d}}\sqrt{\frac{(g_{j,m})^{2}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}}\sigma_{j}^{z}a_{\rm in}^{m},$
@@
$-2\sum_{l\neq j}\sum_{\omega_{c}^{m}\sim\omega_{d}}\left[\frac{g_{j,m}g_{l,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{+}\sigma_{l}^{-}+{\rm h.c.}\right]$
$-2\sum_{\omega_{c}^{m}\sim\omega_{d}}\sqrt{\frac{(g_{j,m})^{2}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}}\left[\sigma_{j}^{+}a_{\rm in}^{m}+{\rm h.c.}\right]$
Finally, we can find a master equation whose effective dynamics for the qubits are equivalent to the above quantum Langevin equations (59)-(60). This is the case for the effective master equation (13) with effective Hamiltonian (14), provided the cavity-mediated decay $\gamma_{\rm cm}^{jl}$, detuning, hopping $J_{\rm cm}^{jl}$, driving $\Omega_{\rm cm}^{j}$, and detuning $\delta_{\rm cm}^{j}$ are given as in Eqs. (39)-(41), with $\omega_{q}^{j}\approx\omega_{d}$. These expressions reduce to Eqs. (15)-(18) in the case of a mono-mode cavity.